OpenStudy (anonymous):

Can someone please check my answers (7 questions)

OpenStudy (anonymous):

1: Solve the equation for y. 4x – 5y = 9 i got y = 4/5x - 9/5

OpenStudy (anonymous):

2: Which ordered pair in the form (a, b) is a solution of this equation? 4a – 5b = 24 A. (–5, –2) B. (6, 0) C. (0, 6) D. (–2, –5) i got (6,0)

OpenStudy (anonymous):

3:Which points have opposite x- and y-coordinates [for example, (2, –2) or (–3, 3)]? (Hint: Some answers may have some, but not all points that name opposites. Choose the answer that contains only points that have opposite x- and y-coordinates for each point named.) A. D, Z B. D, L C. Z, L D. C, D, Z I got B

OpenStudy (igreen):

#1 \(\sf 4x - 5y = 9\) Subtract 4x to both sides: \(\sf -5y = -4x + 9\) Divide -5 to both sides: \(\sf y = \dfrac{4}{5}x - \dfrac{9}{5}\) \(\Huge\color{lime}\checkmark\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

4:Which point is the graph of the ordered pair (2, –4)? A. point A B. point B C. point C D. point D i got point b

OpenStudy (igreen):

#2 \(\sf 4a – 5b = 24\) \(\sf (6, 0)\) Plug in what we know: \(\sf 4(6) – 5(0) = 24\) Multiply: \(\sf 24 – 0 = 24\) \(\sf 24 = 24\) \(\Huge\color{lime}\checkmark\)

OpenStudy (anonymous):

5: In which quadrant does the point (–10, –9) lie? A. Quadrant I B. Quadrant II C. Quadrant III D. Quadrant IV I got B 6:Which answer shows this equation written in standard form? 9 – 4(x – y) = –6x + 6 A. 2x + 4y = –3 B. 2x + 4y = 3 C. –10x + 4y = –3 D. 2x – 4y = 3 i got A 7:Which set of points lies on the given graph? VHS_ALG_S1_07_L203_L303_LQ1-graphic.gif A. (4, 11), (–2, 7), (1, –1) B. (4, 11), (2, –7), (1, 1) C. (–4, 11), (–2, 7), (1, –1) D. (–4, 11), (–2, 7), (1, 1) i got d

OpenStudy (igreen):

And I have to go now.. D:

OpenStudy (igreen):

#5 is correct..

OpenStudy (igreen):

#6 is correct

OpenStudy (igreen):

#7 is correct

OpenStudy (igreen):

#4 is also correct

OpenStudy (anonymous):

ty so much :)

OpenStudy (igreen):

Wait..#3 is correct..lol. My bad.

OpenStudy (anonymous):

it's okay xD